# Gravitational Wave Memory of Primordial Black Hole Mergers

### v1

@@ III.2 Number of detectable events
We consider binaries in any phase of their inspiral. At the time $t_{0}$, the comoving number density of binaries that would merge in the interval $[t_{0}+\tau,t_{0}+\tau+d\tau]$ is given by ${\cal R}(t_{0}+\tau,d){\rm d}\tau$. Assuming GW emission has sufficiently circularized their orbits, i.e. eccentricity is negligible,  [^9] This is a good approximation for binaries formed in the early universe, as GW driven binary evolution is expected to have circularised the orbit well before $t_{0}$ Franciolini:2021xbq . the frequency of GWs emitted by a binary with a time to coalescence $\tau$ is given by Eq. (II.3). Then, the number density of sources at a time $t$ emitting at a frequency $f$ can be calculated as
@@
In Figure 4, we show the number density of PBH binaries as defined in Eq. (22) for representative values of both $f_{\rm PBH}$ and $m_{\rm PBH}$. The decrease as a function of frequency reflects the frequency evolution for fixed PBH masses, which speeds up at higher frequencies following the characteristic ‘chirp’ signature. Consequently, at higher frequencies, a given binary spends less time per frequency bin, and thus the number density at a given frequency decreases. For nearly monochromatic sources, such a steep decline is mitigated by the fact that the PBH population of binaries peaks at merger timescales much larger than the age of the universe.
@@
×$\times\Theta(\hat{\rho}(f_{\rm in},z)-\hat{\rho}_{\rm th}),$ (23)
where $V_{c}$ is the comoving volume and $\hat{\rho}$ identifies the SNR as defined in Eq. (16), applied to memory or the inspiral signal, respectively. The Heaviside step function in Eq. (III.2) introduces the selection bias based on which binary system would have SNR is bigger than the threshold, customary set to $\hat{\rho}_{\rm th}=8$. This, in practice, corresponds to prescribing the observable volume, or equivalently the maximum redshift, up to which binaries with mass $m$ inspiraling at $f_{\rm in}$ would be detectable.
@@ IV Results
We are now in the position to compare the reach of future detectors when searching for PBH binaries through inspiral or memory GW signatures. We consider different future GW experiments: Advanced LIGO (AdLIGO) KAGRA:2013rdx , Einstein Telescope (ET) Hild:2008ng ; Punturo:2010zz ; Hild:2010id , Laser interferometer space antenna (LISA) LISA:2017pwj ; LISACosmologyWorkingGroup:2022jok , and the proposal of HFGW Magnetic Weber Bars (MWB) Domcke:2024mfu . We give more details on the sensitivity of these detectors in App. C.
@@
To understand the structure of these curves in more detail, let us consider for example the sensitivity of LISA to the primary GW signal (solid cyan curve). At small $f_{\rm PBH}$ the merger rate scales as ${\cal R}\propto f_{\rm PBH}^{2}$, after incorporating all relevant suppression factors. We then obtain the bound by imposing $N_{\rm det}\sim 1\sim{\cal R}\times V$, where $V$ is the integrated cosmological volume within which binaries can be observed. For large masses, the SNR is dominated by the contribution near the ISCO frequency, which scales as $f\sim{1}/{m_{\rm PBH}}$, and below the peak sensitivity the SNR integral behaves as $(f^{-7/3}/f^{-6})^{1/2}$, leading to a strong suppression as a function of mass.  [^10] Here, the $f^{-6}$ scaling in the denominator, as well as the corresponding $f^{2}$ factor in the following expression reflects the expected scaling of the LISA sensitivity curve, see App. C. These values are detector-specific, but the sensitivity curves can always be parameterized by a red- and blue-tilted power low in the low and high-frequency regimes, respectively. However, for masses such that the ISCO frequency lies above LISA’s peak sensitivity (roughly $m_{\rm PBH}\lesssim 10^{7}\,M_{\odot}$), the SNR integral change scaling and goes like $(f^{-7/3}/f^{2})^{1/2}$. This modification, together with the effect of the growing cosmological volume at the redshift corresponding to the maximum observable distance, results in an overall bound scaling as $f_{\rm PBH}\sim m_{\rm PBH}^{1/2}$. At masses below approximately $10^{3}\,M_{\odot}$, the duration of the signal in band exceeds the observation time, which reduces the SNR by limiting the number of observable frequency bins. In the small mass limit, primary signals evolve very slowly in frequency and become nearly monochromatic sources, for which the SNR simplifies to $\rho^{2}\sim\tilde{h}^{2}(f_{\rm in})/S_{h}(f_{\rm in})$ Robson:2018ifk . The change of scaling in all constraints around $f_{\rm PBH}\sim 10^{12}$ follows a change in the scaling of the merger rate as a function of $f_{\rm PBH}$. 111111 We iterate that this extrapolation to $f_{\text{PBH}}\gg 1$ is unphysical and we are only interested in a relative comparison between the constraint from inspiral+merger to the one from GW memory: in the low-mass limit, the memory signal can exceed the primary signal. If more sensitive detectors, admittedly quite futuristic, were to become available, all these lines would shift down into the physically meaningful parameter space.
@@
In this work we have focused on the prediction of single loud memory events, but as for the dominant GW signal, the superposition of unresolved events will generate a stochastic background. It has been shown in Refs. Allen:2019hnd ; Zhao:2021zlr that the energy density of such a background decreases faster at low frequencies, $\Omega^{\text{mem}}_{\text{GW}}\propto f$, than for binaries in the inspiral, $\Omega^{\text{ins}}_{\text{GW}}\propto f^{2/3}$. So we expect that for the EU formation scenario the memory background is always subdominant because the cutoff frequency corresponding to the largest binary separation is extremely low, but we leave a detailed discussion to future work.
@@ Appendix A Gravitational wave memory derivation
where $n_{\mu}=-\nabla_{\mu}t+\nabla_{\mu}d$. Far from the source, the GW flux is purely radially outward and only depends on the retarded time $u=t-d$ and on the angular coordinates with respect to the source $\Omega$. Note that the triangular brackets in Eq. (27) are needed for the proper gauge-invariant definition of the GW energy density (see Sec. (1.4) of maggiore2008gravitational ). By defining, as usual, $\bar{h}=h-\eta_{\mu\nu}h^{\alpha}_{\alpha}/2$, in the harmonic gauge where $\partial^{\mu}\bar{h}_{\mu\nu}=0$ for the memory as well, Eq.(26) becomes
@@
It is common in the literature to expand both the main and the memory contribution into spin-weighted spherical harmonics (SWSH) of spin-weight $s=-2$
@@
×$\times\sqrt{\frac{(l-2)!}{(l+2)!}}\begin{pmatrix}l^{\prime}&l^{\prime\prime}&l\\ m^{\prime}&-m^{\prime\prime}&-m\end{pmatrix}\begin{pmatrix}l^{\prime}&l^{\prime\prime}&l\\ 2&-2&0\end{pmatrix},$
where the big parenthesis represents the Wigner $3-j$ symbols, which in this case are only non-zero if $m=m^{\prime}-m^{\prime\prime}$ and $|l^{\prime}-l^{\prime\prime}|\leq l\leq l^{\prime}+l^{\prime\prime}$. From the selection rule above it becomes evident that for a quasi-circular and non-precessing signal, for which the leading GW modes are the $h_{(2,\pm 2)}$, it is easy to see that the memory is predominantly found in the $m=0$ modes with $2\leq l\leq 4$. Therefore, one finds that the leading memory contributions are
$h^{\rm mem}_{(2,0)}=\frac{d}{7}\sqrt{\frac{5}{6\pi}}\int_{-\infty}^{u}{\rm d}u^{\prime}\langle|h_{(2,2)}|^{2}\rangle=60\sqrt{3}h^{\rm mem}_{(4,0)}\,.$
Using Eq. (30) and the fact that $h_{+}=Re[h]$ and $h_{\times}=Im[h]$, in this coordinate system the memory just affects one polarization, which is conventionally chosen to be the $+$ polarization in the source frame with the $z$-axis pointing in the direction of the observer, and the binary plane intersecting the $x$-$y$-plane at an inclination angle $\iota$ (see e.g. favata_nonlinear_2009 ),
$h^{\rm mem}_{+}=\frac{d}{192\pi}\sin^{2}{\iota}(17+\cos^{2}{\iota})\int_{-\infty}^{u}{\rm d}u^{\prime}\langle|h_{(2,2)}|^{2}\rangle$ (32)
@@ Appendix B Memory signals at interferometers
${\cal T}(\hat{\bm{k}}\cdot\hat{\bm{\ell}},f)=Le^{-2\pi ifL}$ $\left[e^{-i\pi fL(1+\hat{\bm{k}}\cdot\hat{\bm{\ell}})}\text{sinc}[\pi fL(1-\hat{\bm{k}}\cdot\hat{\bm{\ell}})]\right.$
$\left.+e^{i\pi fL(1-\hat{\bm{k}}\cdot\hat{\bm{\ell}})}\text{sinc}[\pi fL(1+\hat{\bm{k}}\cdot\hat{\bm{\ell}})]\right]$ (38)
@@
Figure 6: GW signal in LISA for $\theta=\{0,\pi/6,\pi/2\}$ (blue, yellow, green) and $\phi=0$, compared to the approximate expression Eq. (8) (black, gray). Here ^$\hat{\bm{\ell}}_{12}=\hat{e}_{x}$, ^$\hat{\bm{\ell}}_{13}=(1/2,\sqrt{3/4},0)$. The upper panel is for a simple equilateral interferometer, the lower panel uses TDI variables, $s_{\text{TDI}}(t)=s(t)-s(t-2L)$.
@@ Appendix C Power spectral density sensitivity curves
    ×$\times\left(1+\frac{6}{10}\frac{f^{2}}{f_{\star}^{2}}\right)\ ,$ (44)
    $A_{1}={10}/{3L^{2}}$, $L=2.5$Gm, $f_{\star}=19.09$mHz, while
@@
    ×$\times\left[1+\left(\frac{f}{8\textnormal{mHz}}\right)^{4}\right]\ \textnormal{Hz}^{-1}.$ (45)

### v2

@@ III.2 Number of detectable events
We consider binaries in any phase of their inspiral. At the time $t_{0}$, the comoving number density of binaries that would merge in the interval $t_{0}+\tau,t_{0}+\tau+d\tau]$ is given by ${\cal R}(t_{0}+\tau,d){\rm d}\tau$. Assuming GW emission has sufficiently circularized their orbits, i.e. eccentricity is negligible, [^9] This is a good approximation for binaries formed in the early universe, as GW driven binary evolution is expected to have circularised the orbit well before $t_{0}$ Franciolini:2021xbq . the frequency of GWs emitted by a binary with a time to coalescence $\tau$ is given by Eq. (II.3). Then, the number density of sources at a time $t$ emitting at a frequency $f$ can be calculated as
@@
In Figure˜4, we show the number density of PBH binaries as defined in Eq. (22) for representative values of both $f_{\rm PBH}$ and $m_{\rm PBH}$. The decrease as a function of frequency reflects the frequency evolution for fixed PBH masses, which speeds up at higher frequencies following the characteristic ‘chirp’ signature. Consequently, at higher frequencies, a given binary spends less time per frequency bin, and thus the number density at a given frequency decreases. For nearly monochromatic sources, such a steep decline is mitigated by the fact that the PBH population of binaries peaks at merger timescales much larger than the age of the universe.
@@
$\times\Theta(\hat{\rho}(f_{\rm in},z)-\hat{\rho}_{\rm th}),$ (23)
where $V_{c}$ is the comoving volume and $\hat{\rho}$ identifies the SNR as defined in Eq. (16), applied to memory or the inspiral signal, respectively. The Heaviside step function in Eq. (III.2) introduces the selection bias based on which binary system would have SNR bigger than the threshold, customary set to $\hat{\rho}_{\rm th}=8$. This, in practice, corresponds to prescribing the observable volume, or equivalently the maximum redshift, up to which binaries with mass $m$ inspiraling at $f_{\rm in}$ would be detectable.
@@ IV Results
We are now in the position to compare the reach of future detectors when searching for PBH binaries through inspiral or memory GW signatures. We consider different future GW experiments: Advanced LIGO (AdLIGO) at design sensitivity KAGRA:2013rdx , Einstein Telescope (ET) Hild:2008ng ; Punturo:2010zz ; Hild:2010id , Laser interferometer space antenna (LISA) LISA:2017pwj ; LISACosmologyWorkingGroup:2022jok , and the proposal of HFGW Magnetic Weber Bars (MWB) Domcke:2024mfu . We give more details on the sensitivity of these detectors in App. C.
@@
To understand the structure of these curves in more detail, let us consider for example the sensitivity of LISA to the primary GW signal (solid cyan curve). At small $f_{\rm PBH}$ the merger rate scales as ${\cal R}\propto f_{\rm PBH}^{2}$, after incorporating all relevant suppression factors. We then obtain the bound by imposing $N_{\rm det}\sim 1\sim{\cal R}\times V$, where $V$ is the integrated cosmological volume within which binaries can be observed. For large masses, the SNR is dominated by the contribution near the ISCO frequency, which scales as $f\sim{1}/{m_{\rm PBH}}$, and below the peak sensitivity the SNR integral behaves as $(f^{-7/3}/f^{-6})^{1/2}$, leading to a strong suppression as a function of mass. [^10] Here, the $f^{-6}$ scaling in the denominator, as well as the corresponding $f^{2}$ factor in the following expression reflects the expected scaling of the LISA sensitivity curve, see App. C. These values are detector-specific, but the sensitivity curves can always be parameterized by a red- and blue-tilted power low in the low and high-frequency regimes, respectively. However, for masses such that the ISCO frequency lies above LISA’s peak sensitivity (roughly $m_{\rm PBH}\lesssim 10^{7}\,M_{\odot}$), the SNR integral change scaling and goes like $(f^{-7/3}/f^{2})^{1/2}$. This modification, together with the effect of the growing cosmological volume at the redshift corresponding to the maximum observable distance, results in an overall bound scaling as $f_{\rm PBH}\sim m_{\rm PBH}^{1/2}$. At masses below approximately $10^{3}\,M_{\odot}$, the duration of the signal in band exceeds the observation time, which reduces the SNR by limiting the number of observable frequency bins. In the small mass limit, primary signals evolve very slowly in frequency and become nearly monochromatic sources, for which the SNR simplifies to $\rho^{2}\sim\tilde{h}^{2}(f_{\rm in})/S_{h}(f_{\rm in})$ Robson:2018ifk . The change of scaling in all constraints around $f_{\rm PBH}\sim 10^{12}$ follows a change in the scaling of the merger rate as a function of $f_{\rm PBH}$. [^11] We iterate that this extrapolation to $f_{\text{PBH}}\gg 1$ is unphysical and we are only interested in a relative comparison between the constraint from inspiral+merger to the one from GW memory: in the low-mass limit, the memory signal can exceed the primary signal. If more sensitive detectors, admittedly quite futuristic, were to become available, all these lines would shift down into the physically meaningful parameter space.
@@
In this work we have focused on the prediction of single loud memory events, but as for the dominant GW signal, the superposition of unresolved events will generate a stochastic background. It has been shown in Refs. Allen:2019hnd ; Zhao:2021zlr that the energy density of such a background decreases faster at low frequencies, $\Omega^{\text{mem}}_{\text{GW}}\propto f$, than for binaries in the inspiral, $\Omega^{\text{ins}}_{\text{GW}}\propto f^{2/3}$. So we expect that for the EU formation scenario the memory background is always subdominant because the cutoff frequency corresponding to the largest binary separation is extremely low (see also Boybeyi:2024aax ), but we leave a detailed discussion to future work.
@@ Appendix A Gravitational wave memory derivation
where $n_{\mu}=-\nabla_{\mu}t+\nabla_{\mu}d$. Far from the source, the GW flux is purely radially outward and only depends on the retarded time u=tdu=t-d and on the angular coordinates with respect to the source $\Omega$. Note that the triangular brackets in Eq. (27) are needed for the proper gauge-invariant definition of the GW energy density (see Sec. (1.4) of maggiore2008gravitational ). By defining, as usual, $\bar{h}_{\mu\nu}=h_{\mu\nu}-\eta_{\mu\nu}h^{\alpha}_{\alpha}/2$, in the harmonic gauge where $\partial^{\mu}\bar{h}_{\mu\nu}=0$ for the memory as well, Eq.(26) becomes
@@
It is common in the literature to expand both the main and the memory contribution into spin-weighted spherical harmonics (SWSH) of spin-weight s=2s=-2
@@
$\times\sqrt{\frac{(l-2)!}{(l+2)!}}\begin{pmatrix}l^{\prime}&l^{\prime\prime}&l\\ m^{\prime}&-m^{\prime\prime}&-m\end{pmatrix}\begin{pmatrix}l^{\prime}&l^{\prime\prime}&l\\ 2&-2&0\end{pmatrix}$,
where the big parenthesis represents the Wigner 3j3-j symbols, which in this case are only non-zero if $m=m^{\prime}-m^{\prime\prime}$ and $|l^{\prime}-l^{\prime\prime}|\leq l\leq l^{\prime}+l^{\prime\prime}$. From the selection rule above it becomes evident that for a quasi-circular and non-precessing signal, for which the leading GW modes are the $h_{(2,\pm 2)}$, it is easy to see that the memory is predominantly found in the $m=0$ modes with $2\leq l\leq 4$. Therefore, one finds that the leading memory contributions are
$h^{\rm mem}_{(2,0)}=\frac{d}{7}\sqrt{\frac{5}{6\pi}}\int_{-\infty}^{u}{\rm d}u^{\prime}\langle|\dot{h}_{(2,2)}|^{2}\rangle=60\sqrt{3}h^{\rm mem}_{(4,0)}\,.$
Using Eq. (30) and the fact that $h_{+}=Re[h]$ and $h_{\times}=Im[h]$, in this coordinate system the memory just affects one polarization, which is conventionally chosen to be the ++ polarization in the source frame with the $z$-axis pointing in the direction of the observer, and the binary plane intersecting the $x$-$y$-plane at an inclination angle $\iota$ (see e.g. favata_nonlinear_2009 ),
$h^{\rm mem}_{+}=\frac{d}{192\pi}\sin^{2}{\iota}(17+\cos^{2}{\iota})\int_{-\infty}^{u}{\rm d}u^{\prime}\langle|\dot{h}_{(2,2)}|^{2}\rangle$ (32)
@@ Appendix B Memory signals at interferometers
$\Delta T(t,\hat{\bm{\ell}},{\bm{x}}_{1})=\frac{\hat{\ell}^{a}\hat{\ell}^{b}}{2}e_{ab}^{+}\int_{-\infty}^{\infty}\!df\,{\cal T}(\hat{\bm{k}}\cdot\hat{\bm{\ell}},f)\,\tilde{h}(f)\,e^{2\pi if(t-\hat{\bm{k}}\cdot\bm{x}_{1})}.$ (37)
@@
[…]
@@
Figure 6: GW signal in LISA for $\theta=\{0,\pi/6,\pi/2\}$ (blue, yellow, green) and $\phi=0$, compared to the approximate expression Eq. (8) (black, gray). Here $\hat{\bm{\ell}}_{12}=\hat{e}_{x}$, $\hat{\bm{\ell}}_{13}=(1/2,\sqrt{3/4},0)$. The upper panel is for a simple equilateral interferometer, the lower panel uses TDI variables, $s_{\text{TDI}}(t)=s(t)-s(t-2L)$.
@@ Appendix C Power spectral density sensitivity curves
    $\times\left(1+\frac{6}{10}\frac{f^{2}}{f_{\star}^{2}}\right)\ ,$ (44)
    $A_{1}={10}/{3L^{2}}$, L=2.5L=2.5Gm, $f_{\star}=19.09$mHz, while
@@
    $\times\left[1+\left(\frac{f}{8\textnormal{mHz}}\right)^{4}\right]\ \textnormal{Hz}^{-1}.$ (45)